Focus_icon24

 

CLOUD SERVICE STATUS UPDATES

 

 

17/07/2017 14.53pm

We are aware of a brief outage on our hosted solution this afternoon, we believe the issue now to be resolved, if you are still experiencing any issues please don’t hesitate to contact our support team at support@focus-property.co.uk or by phone on 08712500021.


04/07/2017 17.25pm

Service seems to be responding normally, we are looking into the cause.


04/07/2017 16.50pm

Some users are reporting not being able to log in again. We are looking into the issue now.


04/07/2017 9.43am

Everything is now back up and running and all servers are reporting as fully working, all users are able to log in.


04/07/2017 9.28am

It seems that some users are now able to log in again, but we are testing to make sure that full service is resumed.


04/07/2017 9.00am

Customers are reporting issues logging in to the cloud service and we are actively working with our hosting providers to resolve the issue as soon as possible.


09/04/2017 21.00pm

Normal service has been resumed.


09/04/2017 16.16pm

Users are reporting problems with icons not appearing on the remote web portal. Engineers are looking into and we will keep this page updated as we know more.


03/04/2017 10.37am

The cloud service is now back up and running and all users are able to log in. We apologise for any inconvenience caused.


03/04/2017 10.06am

We are experiencing issues with some customers logging in this morning, our engineers are working hard to resolve this issue and we will keep this page up-to-date with the latest information on the situation. 


07/12/2016 16.41pm

We have just received confirmation that the network issue causing some customers to not be able to access the software has been resolved. We apologise again, for the inconvenience but all services are back up and running and all users can log in to the cloud system.


07/12/2016 12.55pm

We now understand that only a handful of customers are still affected by this issue and their IP addresses are being blocked from accessing our servers by something on the internet. Our hosting company are working hard to find out why and resolve this as soon as possible as soon as we have any more information it will be posted here.


07/12/2016 09.30am

Many customers are reporting that the issues from yesterday have been resolved and that they can access the service again, however some customer still appear to have an issue connecting, We apologise for the inconvenience caused and are working closely with all parties involved to ensure normal service is resumed as soon as possible we will update this page regularly this morning with any information we receive.


06/12/2016 17.08pm

Some customers are still unable to access the web portal this will continue to be investigated by the datacenter engineers overnight and should be back to normal service within the next few hours. We will continue to update this page as we receive updates and will of course report when the issue is completely resolved.


06/12/2016 14.00pm

Our hosting provider and Datacenter engineers are aware of the problem and have people at both ends looking into what is causing the issue and we hope to have this resolved as soon as possible.


06/12/2016 11.36am

We are getting reports of issues with some customers logging in, our hosting providers are aware and looking into the issue and it appears to be only some Internet Service Providers as not all users are affected.


22/11/2016 15.46pm

All servers have been restored and users can successfully log in.


22/11/2016 14.10pm

Some customers have reported issues logging into the remote system this afternoon. We are aware of the issue and are actively trying to resolve it.


12/09/2016 12.00pm

We are pleased to announce that parts of the remote service are back up and running. You should now be able to log in but may experience some intermittency issues as we continue to restore the rest of the servers.


12/09/2016 11.05am

Our hosting company’s most senior engineer is investigating why some parts of the remote service are still not online. They will be supplying us with an update in due course and we will be updating this page with news of improvement to the current situation.


12/09/2016 9.40am

The datacenter has reported a power outage which occurred over the weekend and are due to report to us within the next hour, as to why the service is not up and running at this time. Sorry for any inconvenience this is causing you, we will continue to update this page with the status of the service.


12/09/2016 9.00am

We are aware of an issue with our remote servers and are working with our IT company to resolve as soon as possible. Please keep checking this page for updates.


20/07/2016

A small number of our customers in the areas shown below, may experience a loss of telephone and/or broadband services. We hope to have service restored as quickly as possible and apologise for any inconvenience this may be causing. Services can be affected by a variety of reasons such as damage caused by 3rd parties or cable theft:  

Looe – 01203 (estimated clear date 25/07/2016)

Poplar – 0203, 0207 (estimated clear date 25/07/2016)

Dungiven  – 02877 (estimated clear date 25/07/2016) 

Lesmahagow – 01555 (estimated clear date 22/07/2016)

Tewkesbury – 01684 (estmated clear date 22/07/2016)

Hertford – 01469, 01992 (estmated clear date 22/07/2016)

Velindre – 01559 (estmated clear date 22/07/2016)

Yetminster – 01935 (estmated clear date 22/07/2016)

Roothing – 01279 (estmated clear date 22/07/2016)

Gwalchmai – 01407 (estmated clear date 22/07/2016)

Colaton Raleigh – 01395 (estmated clear date 22/07/2016)

Turnberry – 01655 (estmated clear date 22/07/2016)

Frittenden – 01580 (estmated clear date 22/07/2016)

Farnham – 01252 (estmated clear date 22/07/2016)

Crediton – 01363 (estimated clear date 22/07/2016)  

Southrepps – 01263 (estimated clear date 21/07/2016)

Holme – 01485 (estimated clear date 21/07/2016)

Wing – 01296 (estimated clear date 21/07/2016)

Bwlch – 01874 (estimated clear date 21/07/2016)

Gunnislake – 01822 (estimated clear date 21/07/2016)

Portscatho – 01872 (estimated clear date 21/07/2016)

Carterton – 01993 (estimated clear date 21/07/2016)

Nantwich – 01270 (estimated clear date 21/07/2016)

Boncath – 01239 (estimated clear date 21/07/2016)

Black Torrington – 01409 (estimated clear date 21/07/2016)

Kinross – 01577 (estimated clear date 21/07/2016)

Canon Pyon – 01432 (estimated clear date 21/07/2016)

Abergavenny – 01873 (estimated clear date 21/07/2016)

London – 0203,0207,0208 (estimated clear date 20/07/2016)

Bampton – 01398, 01931 (estimated clear date 20/07/2016)

York – 01904 (estimated clear date 20/07/2016)

Caergwrle – 01978 (estimated clear date 20/07/2016)

West Wratting – 01223 (estimated clear date 20/07/2016)

Fritton- 01493 (estimated clear date 20/07/2016)

Goring- 01491 (estimated clear date 20/07/2016)

Little Gaddesden – 01442 (estimated clear date 20/07/2016)

Ickford – 01844 (estimated clear date 20/07/2016)

Dryslwyn – 01558 (estimated clear date 20/07/2016)

Brookeborough – 02889 (estimated clear date 20/07/2016)

Puddletown – 01305 (estimated clear date 20/07/2016)

Melton Constable – 01263 (estimated clear date 20/07/2016)

St Neots – 01480 (estimated clear date 19/07/2016)

Hoyland – 01226 (estimated clear date 19/07/2016)

Hove – 01273 (estimated clear date 19/07/2016)

Wing – 01296 (estimated clear date 19/07/2016)

Ffestiniog – 01766 (estimated clear date 19/07/2016)

Falfield – 01454 (estimated clear date 19/07/2016)

Barking – 0203, 0208 (estimated clear date 19/07/2016)

Wormelow – 01981 (estimated clear date 19/07/2016)

Vauxhall – 0203, 0207, 0208 (estimated clear date 19/07/2016)

Bursledon – 02380, 02381 (estimated clear date 19/07/2016)

Whitchurch – 01275 (estimated clear date 18/07/2016)

Port Dinorwic – 01248 (estimated clear date 18/07/2016)

East Kilbride – 01355 (estimated clear date 18/07/2016)

Bethesda – 01248 (estimated clear date 18/07/2016)

Moulton Chapel – 01406 (estimated clear date 18/07/2016)

Forest Green – 01306 (estimated clear date 18/07/2016)

Unfortunately this morning a number of our customers were left unable to connect to the Remote Desktop Application. This was caused by a BT Service outage. We have now been informed that the service has been fully restored. We invite you to contact our support team on 08712500021 should you continue to be unable to connect.

We apologize for any inconvenience,

Kind Regards,

Focus Micro Systems

 

24/06/2016 – 14.00

Please see below, a message from our contact at the datacentre about a required reboot to services tonight between 6pm and 7pm.

The Applications hosted in RDWeb are now accessible again however additional work is still required to complete the final solution.
The system will be unavailable between 6pm – 7pm tonight so the remaining fixes can be put in place.
All users need to be logged off by 6pm as the works will be started promptly.

Please find below details of planned IT maintenance.

If you have any questions please feel free to contact me.

Maintenance type:                          Planned
Server affected:                                All Servers hosted in Claranet Data Centre
Period start date/time:                  24/06/2016 6pm
Period end date/time:                   24/06/2016 7pm
Service outage expected:             All Services
Length of service outage:             1 hour
Reason for maintenance:              Bring servers up in correct order and install vmware tools update


State: Major Service Incident

Update 16 – 12:00, Friday 24th June 2016

All Claranet shared service platforms have now been restored and we are continuing to work to resolve outstanding individual customer issues. Please raise any outstanding issues via Claranet’s standard contact methods so that engineers can prioritise the restoration of these services. Once we have formally closed the Major Incident we will undertake a full review and provide an Incident report.

We apologise for the inconvenience caused and thank you for your patience.

Time planned for next update: 14:00, Friday 24th June 2016


State: Major Service Incident

Update 14 – 09:30, Friday 24th June 2016

Following restoration of power and shared services engineers are collating a list of outstanding alerts and are focused on the task of recovering services which continue to be impacted.

We apologise for the inconvenience caused and thank you for your patience.

Time planned for next update: 10:30, Friday 24th June 2016


State: Major Service Incident

Update 13 – 08:15, Friday 24th June 2016

The final VDC Host is being restored and then the main task of recovering customers will start with immediate effect.


State: Major Service Incident

Update 13 – 08:15, Friday 24th June 2016

The final VDC Hosts is be restored and then the main task of recovering customers will start with immediate effect.

Thank you for your patience.

Time planned for next update: 09:30, Friday 24th June 2016


State: Major Service Incident

Update 12 – 07:15, Friday 24th June 2016

The remaining VDC Hosts will be restored with an ETA of 1.5hrs. Claranet engineers continue to work and remain focused on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 08:30, Friday 24th June 2016


State: Major Service Incident

Update 11 – 04:15, Friday 24th June 2016

Hosts continue to be restored slowly. Claranet engineers continue to work through the night and remain focused on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 07:30, Friday 24th June 2016


State: Major Service Incident

Update 9 – 03:15, Friday 24th June 2016

Service continue to be restored slowly, please be assured Claranet continue to focus with the highest priority on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 04:30, Friday 24th June 2016


State: Major Service Incident

Update 9 – 02:15, Friday 24th June 2016

vCentre is back up and we have an ETA of apprx 2 hours to restore VDC.

Claranet continue to focus with the highest priority on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 03:30, Friday 24th June 2016


State: Major Service Incident

Update 8 – 01:15, Friday 24th June 2016

Claranet continue to focus with the highest priority on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 02:30, Friday 24th June 2016


State: Major Service Incident

Update 7 – 00:30, Friday 24th June 2016

GS 2 have confirmed that full power and resilience has been restored. Investigation into root cause continues.

Claranet continue to focus with the highest priority on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 01:30, Friday 24th June 2016


State: Major Service Incident

Update 6 – 00:30, Friday 24th June 2016

GS 2 have confirmed that full power and resilience has been restored. Investigation into root cause continues.

Claranet continue to focus with the highest priority on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 01:30, Friday 24th June 2016


State: Major Service Incident

Update 6 – 23:00, Thursday 23rd June 2016

Claranet are continuing to restore customer services.

Power is restored albeit we are still running on protected mains with no UPS support.

All co-lo customers have power to their racks, if any customers require remote hands assistance please contact the service desk via the normal methods.

We have recovered our Managed Virtual Hosting (MVH) platform with customers servers back online

We continue to focus with the highest priority on restoring service across our Virtual Data Centre platform and those customers with managed services.

Thank you for your patience.

Time planned for next update: 00:30, Friday 24th June 2016


State: Major Service Incident

Latest update:

Update 5 – 22:00, Thursday 23rd June 2016

As advised power has been restored to Global Switch 2 and Claranet have restored power to our suite.

Engineers continue to work to restore service however we have no ETA for full restoration. Internal hourly calls continue to drive progress.

Thank you for your patience.

Time planned for next update: 23:00, Thursday 23rd June 2016


State: Major Service Incident

Service Type/ Products impacted: All Services in Global Switch 2

Incident Summary: Loss of Service

Business Impact: Loss of Service

Latest update:

Update 3 – 19:00, Thursday 23rd June 2016

We have had confirmation that power to Global Switch 2 has now been restored and we are now working on restoring services in a controlled manner. In parallel we are working to understand root cause.

Thank you for your patience.

Time planned for next update: 20:30, Thursday 23rd June 2016